Question 34

  1. Let Q = quantity produced, P = selling price per unit, VC = variable cost per unit, and TFC = total fixed cost. Which of the following equations is correct?

    Profit = Q × (P – VC) + TFC

    a.      Profit = Q × (P – VC) - TFC

    Profit = Q × (P – VC – TFC)

    Profit = Q × (P – VC + TFC)

2.85714 points

Question 35

  1. The Hancock Corp. plans to sell its products for $100 each. Its variable cost per unit = $80, and its fixed costs for the year = $100,000. What level of sales will Hancock need to break even?

    1,000 units

    5,000 units

    10,000 units

    20,000 units
Answer #1

Answers:

1) Profit = Q*(P-VC) -TFC

2) 5,000 Units

Profit Equation = Quanitity/( Selling price - Variable cost per unit ) - Total fixed cost

Profit = Q*( P-VC) -TFC

Break Even Units = Fixed cost / ( Selling price - Varaible cost ) = 100,000 /(100-80) = 5000 Units